OpenStudy (anonymous):

help will give medal Traveling at 55 miles per hour, how many minutes, rounded to the nearest whole number, does it take to drive 310 miles from Boston to Syracuse? (5 points) 6 338 365 3,300

OpenStudy (anonymous):

\[ \frac{ 155 }{ 3 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

still got the same answer

OpenStudy (anonymous):

\[\frac{ 62 }{ 11 }\]

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

\[5\frac{ 7 }{ 11 }\]

OpenStudy (anonymous):

5.63